consider a population of 5-lb sacks of potatoes for which the variable of interest is the number of potatoes,for which the variable of interest is the number of potatoes.It is a discrete uniform distribution on the intergers 12 through 18.

find the mean ,variance ,and standard deviation of this sampling distribution of the mean for n=2


1
Expert's answer
2021-08-12T14:21:13-0400

The sampling distribution of the sample mean is as follows:



Mean of the sample mean,

"E\left(\overline{x}\right)=\sum \:\overline{x}P\left(\overline{x}\right)=15"

"E\left(\overline{x}^2\right)=\sum \:\overline{x}^2P\left(\overline{x}\right)=227"


"Var\left(\overline{x}^2\right)=E\left(\overline{x}^2\right)-\left[E\left(\overline{x}\right)\right]^2"

"Var\left(\overline{x}^2\right)=227-15^2=2"


Standard deviation of "\overline{x}=SD\left(\overline{x}\right)=\sqrt{Var\left(\overline{x}^2\right)}"

"SD\left(\overline{x}\right)=\sqrt{2}\approx 1.414"